Q26.4: The heating element of a hair dryer dissipates 1500 W when connected to a 150 V outlet. You may want to review (Pages 841 - 843) .

What is its resistance? Express your answer to two significant figures and include the appropriate units.
Answer #1

Here,

Power , P = 1500 W

Voltage , V = 150 V

let the resistance is R

as P = V^2/R

R = V^2/P

R = 150^2/1500

R = 15 Ohm

the resistance is 15 Ohm